if the sum of the series 2,5,8,11..... is 60100, then n is equal  to??????

Grade:12th pass

6 Answers

aarthi s
8 years ago
Use the formula S=n/2[2a+(n-1)d]
substitute the values a=2, d=5-3 =2 & S=60100
Ans 200
mansi dabriwal
8 years ago
i think d should be 5-2=3
Shubham Singh
8 years ago
n = 200
it is an AP with common difference 3.
mansi dabriwal
8 years ago
yah i got it...
Sujit Yadav
8 years ago
ans 200
ng29
8 years ago
it is quite easy question 
it is nothing but an increasin ap with   a=2  d=3  and nth term is 60100
using an=a+(n-1)d      an is the nth term
we get n=200
 
approve if useful